OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary.55 mi / h = ____ ft / s

hero (hero):

\[\frac{55 \space\cancel{\text{mi}}}{\cancel{\text{hr}}} \times \frac{\cancel{\text{hr}}}{3600 \space\text{sec}} \times \frac{5280\space\text{ft}}{\cancel{\text{mi}}}=?\]